Sol Summary November 10th

MDRS Daily Summary Report for sol 0.7

Summary Title: Some reflections about the sustainability

Mission Status: I consider the day very productive and it is difficult to know when something will surprise us. At 6 PM I was surprised with a good discussion related with environmental sustainability about how to save water.

Sol Activity Summary: In the EVA today the 2 ATVs were used.

Look Ahead Plan: Tomorrow, We are planning an EVA using the ATV in the morning to collect soil and plants samples. Julio and Carmen will develop their research in the Hab.

Anomalies in work: Nope.

Weather: 9o C, 49o F

Crew Physical Status: Fine.

EVA: today happened EVA with the ATV in the Tank Wash collecting vegetation samples.

Reports to be file: We sent in CapComm.

Support Requested: lights and fire alarms for the lab; GPS for EVAs and new radios.

Science Report November 10th

Science Report
Camila Castillo
10 Novermber 2017

Mars is pretty nice this night. There’s no cold in our little and warm hab and the night doesn’t seem as cold as ever.

This Science Report should be short because the main reason I’m writing this is because of the ovens at the lab. I’m using a little one and I would like to know if there is a problem with that one. I tried to set up another temperature and it just keeps getting warmer inside and doesn’t really mantain the temperature I originally intended. There are other two ovens, but it seems that they are a lot warmer than I need. I just want to set them up at a temperature near 25 °C.

As for the experiments, they are going well. It seems that the PC plates were well made and we could manage to keep an sterilized condition for making them. I’m expecting a package of Nutrient broth and Nutrient agar and it should be arriving this evening or even tomorrow. It will help a lot for my microbiological experiments.

Today we collected some soil and plant samples on the EVA. Our crew geologist keeps working on their characterization and I could train a bit of my plant systematics. I’m not an expert botanist, but I suppose you need to know a little bit of everything when you’re a scientist on Mars! (I’m learning so much about soils, it’s just crazy). It would be interesting if we get to know the microorganisms that live with these plants. This would be possible with the help of the Nutrient agar (That’s the one we usually use at my lab for this kind of samples).

I have some pretty crazy results on my pilot on the effects of dishwater on my raddish seeds. The ones that are growing the most are the seeds with 10% water and 90% dishwater! It’s really bugging me. I’ll need to follow them to know if they have some kind of alteration during their growing.

Operations Report November 8th

Crew 182 Operations Report Sign-On 08Nov2017

SOL: 6

Carmen Atauconcha

Non-nominal systems: Working

Generator: On

Solar: For now, not given power to the Hab, Used until 5pm approximately.

Diesel : 60%

Propane – 51psi

Gasoline (5 Gallon containers for ATV) – 0 gallons

Water (trailer) – empty

Water (static) – 30 cm from top

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 760 gallons

ATV’s Used: No

Oil Checked: motos y carros: No

Atv # Fuel Used Gals: 1/4 gallons

Tires Status: Perfect

# Hours Used Day: 0 hours

Phobos used: No

Hours:0

Beginning charge: –

Ending charge: –

Currently charging: –

Deimos used: No

Hours: 0hrs

Beginning charge: –

Ending charge: –

Currently charging: –

HabCar used?: No

Notes and Comments:
Today, we started to use the solar panels. We monitored the state of charging of them every hour. The collected data is down:

TIME % CHARGE OF THE SOLAR PANELS
8.43 am 98%
9.13 am 98%
9.43 am 98%
10.09 am 100%
10.39 am 100%
12.20 pm 100%
13.01 pm 98%
13.38 pm 96%
13.46 pm 95%
13.56 pm 94%
14.11 pm 93%
14.34 pm 100%
15.19 pm 100%
16.28 pm 98%
17.31 pm 93%

Also today we had a walking EVA, and tomorrow we are going to have an EVA session. I need to let you know that the gasoline will be available until tomorrow. Therefore, for the next EVA’s we are going to need gasoline.

Greenhab Report November 9th

Green Hab Oficer Atila Meszaros

Sol 6 – 09/11/2017

Ambient (no heating or/cooling): none

Ambient with window/door open: door opened at 08:32; closed at 18:00

Shade cloth on/off: off

Average temperature:

Time *08:32 10:35 *13:55 17:30
Temperature inside (°C) 22 36 34 16
Temperature outside (°C) 8
% Humidity 16 16 16 16

*Time when seed of a former experiment where watered. (There are tomatoes, melons, peppers, cucumbers and beans)

Changes to crops: None

Daily water usage for crops: 2 ½ galons

Time(s) of watering for crops: 08:32 and 13:55

Changes to research plants: none

Daily watering and amount of water used: 1 L approximately

Narrative:

Hi, Graeme! Atila in here, yes, like the Hun. I hope you are having an amazing night.

Sol 6 was a very short day at the Green Hab (GH) because we had the visit of Globo Tv, as you should know by now. I started to rearranged everything at first hour, it’s a huge mess. There some many materials! Although, most of them are useless, there are some items that can be very valuable for next crews, but I’ll need more help tomorrow if I want to finish what I am aiming.

Respecting my project. Half of the seeds remain at the GH; the other half is in the growing chamber. For all the seed of the experiment, I’m using less than a liter of distilled water per day.

The project wants to prove that Peruvian native crops can, indeed, germinate and grow on Mars soil. I’m working with quinoa, amaranthus and mustard (as a control).

EVA Report November 9th

EVA Report Sol 6

The EVA started at 10:50 AM today. Knowing that Brazilian TV was coming, I didn’t know for sure when they’d be here and ready to film us outside. They also needed to interview a few of the crewmembers on their projects before we could start. They wanted some footage of us on the Rover and ATVs, which we drove about 100 meters past the MDRS sign and back to the Hab area. I understand that I did not request the use of the rovers for today’s EVA, but I didn’t know they’d want footage of us on the rovers until this morning. Something to note was that the red ATV had some trouble starting compared to yesterday. I had to pump the throttle in order to get it going.

After parking the rovers, we collected our samples from Sample Location #1, the first location we sampled from on EVA #1. Once we got the sample, we walked into the small gully just north of the Hab. We took samples from 3 different locations along the hillside west of the hab. The “popcorn” like surface of the regolith can be rather unstable, as I found out the hard way (by falling). Nothing broke, I wasn’t injured, the helmet was intact, and the fan was still working, so luckily I survived! After photographing the area, we returned to the hab and recompressed by 12:45 PM.

In Hindsight, I’ll have to be more careful on even very shallow slopes, as it doesn’t take much too loose traction, even when wearing boots. To make sure no one else falls by making the same mistake, I’ll have to brief the crewmembers on how to lower their center of gravity (by squatting) when ascending or descending hills. I did not have time to brief the crew members on radio Comms voice procedures since we had the Brazilian TV crew here today. This will have to be done before tomorrow’s EVA.

Commander Report November 9th

My soundtrack of last night: ACDC live in River Plate stadium, Buenos Aires.:) Camila did a delicious fried rice. Thank You!:)

4h30m – I woke up early today. Sometimes the heating system switches off. So we need to check. This is also a possibility in Mars. Many times, in Mars, the heating systems will not function properly. So, will be necessary a lot of resilience about. Try to fix or deal with the consequences.

The commander’s room is unique with natural lighting. This helps me to woke up early.

7h02m – I walked 40 minutes, something around 2,5 km. Was 1 degreeC. In the beginning of the exercise was very difficult. I never did exercises in a temperature like this. This is also about resilience. In Mars the temperature can be so low. According to Space.com, on average, the temperature on Mars is about minus 80 degrees F (minus 60 degrees C). In winter, near the poles temperatures can get down to minus 195 degrees F (minus 125 degrees C). A summer day on Mars may get up to 70 degrees F (20 degrees C) near the equator, but at night the temperature can plummet to about minus 100 degrees F (minus 73 C). (Source: https://www.space.com/16907-what-is-the-temperature-of-mars.html)

Something very pleasant today was watch the sunrise in the mountains. I perceive how We are not aware about how the cosmic influences are affecting life. Just look the sun! It is a star and it is so close and far from us, influencing so many cycles of our life, from the agriculture to our communication systems.

8h40m – The Brazilian TV Globo arrived earlier than We imagined. From the MDRS corridors I talked with Luigi Sofio (reporter) and Lucio Rodrigues (cinegraphist) and introduced myself. Was good to talk with someone on portuguese. Since their arrival until 13h, We presented all the attention to Them. They did images of our systems and procedures and also the research activities.

11h – We did the EVA activity. Was very hard. Was my first EVA. Would possible to see how the spacesuit would restricts simple human activities. I felt excited with it. This moves me. A spacesuit would be my future?

We did a lot of work today, outside. Today was very shiny and I considered that Brazilian TV did very, very good images. What sincronicity We have a bright day as today.

In the EVA We collected soil samples and visit the ravines and some mountains of Mars. Was hard climb some mountains wearing the spacesuits, considering the limitations of movement. It is necessary presents a great attention and care to avoid an accident. But the EVA was a challenger and an achievement.

13h – I prepared the lunch: small bakes of rice and cheddar with broccoli and peas.

13h40m – We lunched. We had a lot to clean up.

15h40m – The guys of the EVA: Me (Julio), Brandon and Atila felt tired, because the sun radiation was to high today. So everybody decide take a nap.

17h40m – I woke up from my nap. It is good remember about the CapComm later. Good time to listen Slayer and finish the reports. J

18h20m – Camila and Carmen prepared sweet rice for tonight.

19h – CapComm. Today I walked the total of 5,8 km.

Sol Summary November 9th

MDRS Daily Summary Report for sol 0.6

Summary Title: A shine and hot day.

Mission Status: The internet is so low today. The day circles around Globo TV visit. They arrived in MDRS at 8h40m and leaved 13h. We support them in interviews, images and guidelines.

Sol Activity Summary: In the EVA today We walked more, compared to other day and also was so shiny.

Look Ahead Plan: Tomorrow, We are planning an EVA in the morning to study some plants around MDRS to the research of Camila. Julio Rezende will work in his book. Atila and Carmen will develop their research also.

Anomalies in work: Nope.

Weather: 9o C, 49o F

Crew Physical Status: Fine.

EVA: today happened EVA with the ATV hide and walking in the ravines and Mars mountains.

Reports to be file: We sent in CapComm.

Support Requested: None.

Science Report November 9th

Mars has been also great today! We had some sunshine, but it wasn’t as warm as other days. In fact, I had problems finding some sunshine for my seeds. But alright, continuing with the Science report:

The autoclave is already set and working! We read carefully the manual to use this, because I’m particularlly fearful of that machine. But I noticed that it requires a great amount of destillated water for running. I would like to ask about this issue because I had to use a gallon and a half for one sterelization cycle. I’ll need to do another one in the next week, so I would like to know if it is possible to get more gallons of destillated water.

We also need to figure out the position of the machines at the lab. The autoclave requires a good table, so it occupies the only lab table at the lab. I will not be using the hood for my project, so I need to move it somewhere else. Right now, the lab is pretty messy.

As for the projects, there has been some soil sampling today during the EVAs. These samples have been collected for both typing and using during the germination experiments.

No photos are going to be attached to this report, because my main concern is with the destillated water.